# Novel topological subclass in Hor̆ava-Lifshitz black holes

### v1

@@ I Introduction
The field of black hole thermodynamics has long attracted significant attention, offering essential insights into the fundamental links between quantum information, statistical mechanics, and geometric properties qq1 ; qq2 . Phase transitions are widespread in thermodynamic systems. By taking AdS space as the background and studying Schwarzschild black holes as the thermodynamic system, it has been discovered that the critical temperature of a black hole determines its stability, a process known as the Hawking-Page phase transition ch01 , which has major consequences for gauge field theory, particularly through the AdS/CFT duality. The fundamental principle behind the comprehensive investigation of black hole chemistry is to view the cosmological constant as a thermodynamic quantity (pressure), whose conjugate variable is naturally associated with the black hole volume. This approach reveals the novel phase transitions and intricate phase structures qq3 ; qq4 ; qq5 ; qq6 .
@@
Despite notable advances in black hole thermodynamics in recent times, comprehending their universal properties remains a difficult endeavor. By leveraging Duan’s theory of $\varphi$-mapping topological currents qq6.1 ; qq6.2 , topological methods offer a fresh perspective. Wei et al. treat black hole states (solutions) as topological defects, and various black hole systems are sorted into three categories, each determined by the topological numbers linked to their respective topological charges qq7 . They also introduced the topological method for studying black hole critical points, assigning each critical point a topological charge through a temperature-dependent function, they assigned topological charges to each critical point, classifying them into two types: conventional critical points and newly identified critical points qq9 . Following this, by investigating the asymptotic behavior of the vectors, they introduced a framework that categorized black hole states into four universal topological groups, providing a clear system for classifying stable and unstable black hole configurations as the radius of the event horizon increases qq8 . For further examples of recent developments, see references ws1 ; ws2 ; ws3 ; ws4 ; ws5 ; ws6 ; ws7 ; ws7.1 ; ws7.2 ; ws7.3 ; ws7.4 ; ws7.5 ; ws8 ; ws9 ; ws9.1 ; ws10 ; qq11 ; ri1 ; qq12 ; qq13 ; wd1 ; wd2 ; wd3 ; wd4 ; wd5 ; wd6 ; wd7 ; kj1 ; kj2 ; kj3 ; wd8 ; wd9 ; wd10 ; xx1 ; xx2 ; xx3 ; xx4 . Building on this, Wu et al. found that black holes in the framework of gauge supergravity theory qq15 ; qq15.1 , and multiply rotating Kerr-AdS black holes qq14 exhibit an unusual thermodynamic stability at the low-to-high temperature limits of the Hawking temperature, which cannot be accounted for by the four existing topological classifications. This led to the expansion of the classification including five topological classes and three subclasses.
@@
where, the inverse temperature parameter, denoted as $\tau$ or equivalently $\beta$, corresponds to the reciprocal of the Hawking temperature ($\tau=1/T$), as detailed in qq8 . With this relation, Equation (1) reduces to
@@
Next, we analyze the asymptotic characteristics of the vector fields near the boundary, as outlined in Eqs. (9)–(13); the boundary in this context is defined by the contour $C=I_{1}\cup I_{2}\cup I_{3}\cup I_{4}$, where
@@ III The newly defined topological subclass: $\ddot{W}^{1-}$
Figure 2: On the $r_{h}-\beta$ plane, the zero points of the vector $\phi^{r_{h}}$ for the charged HL black hole, using parameters $k/r_{0}=0$, $q/r_{0}=1$, and $Pr_{0}^{2}=0.01$. The stable branch is represented by the red curve.
Figure 3: On the $r_{h}-\beta$ plane, the zero points of the vector $\phi^{r_{h}}$ for the charged HL black hole, with parameters $k/r_{0}=-1$, $q/r_{0}=1$, and $Pr_{0}^{2}=0.01$. The blue unstable branch (w=1w=-1) and the red stable branch ($w=+1$) meet at the annihilation point (AP) marked by the pink dot. The count of stable and unstable states yields two stable and one unstable state, leading to a total topological number of W=1+1+1=1W=-1+1+1=1.
The thermodynamic quantities, including temperature, entropy, and electromagnetic potential are given by
@@
Figure 4: On the $r_{h}-\beta$ plane, the zero points of the vector $\phi^{r_{h}}$ for the charged HL black hole, with the parameters $k/r_{0}=1$, $q/r_{0}=1$, and $Pr_{0}^{2}=0.01$. The stable branch is represented by the red curve.
Figure 5: On the $r_{h}-\beta$ plane, the zero points of the vector $\phi^{r_{h}}$ for the charged HL black hole, with parameters $k/r_{0}=0$, $q/r_{0}=1$, and $Pr_{0}^{2}=0.01$. At the intersection of the blue unstable (w=1w=-1) and red stable ($w=+1$) branches, denoted by the pink annihilation point (AP).
@@ III.1 Black holes in the context of the canonical ensemble
In this subsection, we will examine the the asymptotic temperature behavior of charged black holes with different horizons (spherical, flat and hyperbolic) in the context of the canonical ensemble, which enables the topological classification of black holes. Under this ensemble, the volume, Hawking temperature, and particle number are held constant. By inserting equation (20) and equation (23) into equation (1), we can derive
@@
Figure 6: On the $r_{h}-\beta$ plane, the zero points of the vector $\phi^{r_{h}}$ for the charged HL black hole, with parameters $k/r_{0}=-1$, $q/r_{0}=1$, and $Pr_{0}^{2}=0.01$. The annihilation point (AP) is marked by a pink dot, while the generation point (GP) is indicated by a black dot. Under the condition $\beta<\beta_{c1}=9.61788$, the system exhibits two unstable small black hole states and one stable large black hole state, resulting in the total topological number W=11+1=1W=-1-1+1=-1. For the temperature range $\beta_{c2}=23.2166<\beta<\beta_{c3}=23.2196$, the system consists of two unstable states and one stable small black hole, yielding the total topological number W=1+11=1W=-1+1-1=-1.
@@ III.2 Black holes in the context of the grand canonical ensemble
In this subsection, we examine the asymptotic temperature behavior of black holes with various types of horizons (spherical, flat, and hyperbolic) under the grand canonical ensemble. Here, the system can exchange both energy and charge with the external environment, while its temperature, volume, and chemical potential are constrained. Within this framework, the generalized free energy $(\mathcal{F})$ is defined as
@@
In Figure (6), two independent curves are shown. In the first curve located in the upper left area, two branches of black holes are observed. As the event horizon radius expands, the black holes undergo a sequence of unstable $(\omega=-1)$ and stable $(\omega=1)$ phases. The two states meet and ultimately vanish at the critical point, referred to as the annihilation point, where $\beta_{c1}=9.61788$. The second curve reveals that within a particular temperature range, here are three branches of small black holes, with their stability alternating between unstable $(\omega=-1)$, stable $(\omega=1)$, and unstable $(\omega=-1)$ states as the radius of the event horizon increases. When the inverse temperature parameter satisfies $\beta<\beta_{c1}=9.61788$, three black hole states coexist, leading to a total topological number of W=11+1=1W=-1-1+1=-1. The system’s asymptotic behavior differs from the eight previously established thermodynamic topological structures. In this case, the inner region of the black hole is unstable, whereas the outer region remains stable. In the low-temperature limit, the system features a single unstable small black hole state. In the high-temperature limit, the system supports two unstable small black hole states along with one stable large black hole state. As the inverse temperature increases, the system shifts to a state with only one unstable small black hole. For temperatures within the range $\beta_{c2}=23.2166<\beta<\beta_{c3}=23.2196$, the system contains two unstable small black hole branches and one stable large black hole branch, yielding a total topological number of W=1W=-1. Notably, throughout all values $(\beta)$, the topological number remains consistently 1-1, and the system’s topological classification matches that of the $W^{1-}$ classes. As a result, we suggest that this system belongs to a new subclass, $\ddot{W}^{1-}$, within the topological classification framework.
@@
In this analysis, we explore the asymptotic thermodynamic characteristics of the newly discovered topological subclass. Initially, the system consists of at least three black hole states: with an increasing event horizon radius, the system transitions through three phases: two unstable states $(\omega=-1)$ and one stable state $(\omega=1)$. In these phases, the heat capacities are negative for the first two states and positive for the final, stable state. Subsequent states appear in pairs, where each pair includes one state with a winding number of 1-1 and one with +1+1. The arrangement of winding numbers does not follow a simple alternating pattern but instead begins with [,+][-,+]. This ensures that the innermost black hole is associated with $\omega=-1$ (an unstable state), and the outermost one with $\omega=+1$ (a stable state), forming a topological classification represented by [,+][-,+].
@@ Acknowledgments
We are grateful to Dr. Di Wu for useful discussions on the topological subclass. This work is supported by the National Natural Science Foundation of China (NSFC) under Grants No. 12265007, 12405081, 11865018. by the Doctoral Foundation of Zunyi Normal University of China (BS [2022] 07).

### v2

@@ I Introduction
The field of black hole thermodynamics has long attracted significant attention, offering essential insights into the fundamental links between quantum information, statistical mechanics, and geometric properties qq1 ; qq2 . Phase transitions are widespread in thermodynamic systems. By taking AdS space as the background and studying Schwarzschild black holes as the thermodynamic system, it has been discovered that the critical temperature of a black hole determines its stability, a process known as the Hawking-Page phase transition ch01 , which has major consequences for gauge field theory, particularly through the AdS/CFT duality. The fundamental principle behind the comprehensive investigation of black hole chemistry is to view the cosmological constant as a thermodynamic quantity (pressure), whose conjugate variable is naturally associated with the black hole volume. This approach reveals the novel phase transitions and intricate phase structures qq3 ; qq4 ; qq5 ; qq6 ; ka1 .
@@
Despite notable advances in black hole thermodynamics in recent times, comprehending their universal properties remains a difficult endeavor. By leveraging Duan’s theory of $\varphi$-mapping topological currents qq6.1 ; qq6.2 , topological methods offer a fresh perspective. Wei et al. treat black hole states (solutions) as topological defects, and various black hole systems are sorted into three categories, each determined by the topological numbers linked to their respective topological charges qq7 . Based on this, various types of black hole solutions have been studied as topological defects fm1 ; fm2 ; fm3 ; fm4 ; fm5 ; fm6 ; fm7 ; fm9 ; fm10 ; fm11 ; fm12 ; fm13 ; fm14 ; fm15 ; fm16 ; fm17 ; fm18 ; fm19 ; fm20 ; fm21 . They also introduced the topological method for studying black hole critical points, assigning each critical point a topological charge through a temperature-dependent function, they assigned topological charges to each critical point, classifying them into two types: the conventional critical point and the novel critical point qq9 . Recent studies have widely investigated the topological classification of critical points across different black holes in gravitational backgrounds tp1 ; tp2 ; tp3 ; tp4 ; tp5 ; tp6 ; tp7 ; tp8 ; tp9 ; tp10 . Following this, by investigating the asymptotic behavior of the vectors, they introduced a framework that categorized black hole states into four universal classifications, providing a clear system for classifying stable and unstable black hole configurations as the radius of the event horizon increases qq8 . For further examples of recent developments, see references uf1 ; uf2 ; uf3 ; uf4 ; uf5 . Building on this, Wu et al. found that black holes in the framework of gauge supergravity theory qq15 ; qq15.1 , and multiply rotating Kerr-AdS black holes qq14 exhibit an unusual thermodynamic stability at the low-to-high temperature limits of the Hawking temperature, which cannot be accounted for by the four existing topological classifications. This led to the expansion of the classification including five topological classes and three subclasses. In addition, topological methods have also been employed in the study of light rings lr1 ; lr2 ; lr3 ; lr4 and timelike circular orbits lr5 ; ws1 ; ws3 .
@@
where, $\tau$ denotes the inverse temperature parameter, the free energy reduces to an on-shell quantity only when $\tau=\beta=1/T$, as detailed in qq8 . With this relation, Equation (1) reduces to
@@
Next, we analyze the asymptotic behavior of the vector fields near the boundary, as outlined in Eqs. (9)–(13); the boundary in this context is defined by the contour $C=I_{1}\cup I_{2}\cup I_{3}\cup I_{4}$, where
@@ III The newly defined topological subclass: $\ddot{W}^{1-}$
Figure 2: In the $r_{h}-\beta$ plane, the zero points of the vector $\phi^{r_{h}}$ for the charged HL black hole, using parameters $k/r_{0}=0$, $q/r_{0}=1$, and $Pr_{0}^{2}=0.01$. The stable branch is represented by the red curve.
Figure 3: In the $r_{h}-\beta$ plane, the zero points of the vector $\phi^{r_{h}}$ for the charged HL black hole, with parameters $k/r_{0}=-1$, $q/r_{0}=1$, and $Pr_{0}^{2}=0.01$. The blue unstable branch (w=1w=-1) and the red stable branch ($w=+1$) meet at the annihilation point (AP) marked by the pink dot. The count of stable and unstable states yields two stable and one unstable state, leading to a total topological number of W=1+1+1=1W=-1+1+1=1.
The thermodynamic quantities, including the temperature, entropy, and electromagnetic potential are given by
@@
Figure 5: In the $r_{h}-\beta$ plane, the zero points of the vector $\phi^{r_{h}}$ for the charged HL black hole, with parameters $k/r_{0}=0$, $q/r_{0}=1$, and $Pr_{0}^{2}=0.01$. At the intersection of the blue unstable (w=1w=-1) and red stable ($w=+1$) branches, denoted by the pink annihilation point (AP).
@@ III.1 Black holes in the context of the canonical ensemble
In this subsection, we will examine the the asymptotic behavior of charged black holes with different horizons (spherical, flat and hyperbolic) in the context of the canonical ensemble, which enables the topological classification of black holes. Under this ensemble, the volume, Hawking temperature, and particle number are held constant. By inserting equation (20) and equation (23) into equation (1), we can derive
@@
Figure 6: In the $r_{h}-\beta$ plane, the zero points of the vector $\phi^{r_{h}}$ for the charged HL black hole, with parameters $k/r_{0}=-1$, $q/r_{0}=1$, and $Pr_{0}^{2}=0.01$. The annihilation point (AP) is marked by a pink dot, while the generation point (GP) is indicated by a black dot. Under the condition $\beta<\beta_{c1}=9.61788$, the system exhibits two unstable small black hole states and one stable large black hole state, resulting in the total topological number W=11+1=1W=-1-1+1=-1. For the temperature range $\beta_{c2}=23.2166<\beta<\beta_{c3}=23.2196$, the system consists of two unstable states and one stable small black hole, yielding the total topological number W=1+11=1W=-1+1-1=-1.
@@ III.2 Black holes in the context of the grand canonical ensemble
In this subsection, we examine the asymptotic behavior of black holes with various types of horizons (spherical, flat, and hyperbolic) under the grand canonical ensemble. Here, the system can exchange both energy and charge with the external environment, while its temperature, volume, and chemical potential are constrained. Within this framework, the generalized free energy $(\mathcal{F})$ is defined as
@@
In Figure (6), two independent curves are shown. In the first curve located in the upper left area, two branches of black holes are observed. We find that the asymptotic behavior of the inverse temperature parameter coincides with that of the $W^{1-}$ class, which is described as follows
$\beta\left(r_{m}\right)=0,\quad\beta(\infty)=\infty.$ (33)
@@
It can also be found that with the increase of the horizon radius, the black holes undergo a sequence of unstable $(\omega=-1)$ and stable $(\omega=1)$ phases. The two states meet and ultimately vanish at the critical point, referred to as the annihilation point, where $\beta_{c1}=9.61788$. The second curve reveals that within a particular temperature range, here are three branches of small black holes, with their stability alternating between unstable $(\omega=-1)$, stable $(\omega=1)$, and unstable $(\omega=-1)$ states as the radius of the event horizon increases. When the inverse temperature parameter satisfies $\beta<\beta_{c1}=9.61788$, three black hole states coexist, leading to a total topological number of W=11+1=1W=-1-1+1=-1. The system’s asymptotic behavior differs from the eight previously established thermodynamic topological structures. In this case, the inner region of the black hole is unstable, whereas the outer region remains stable. In the low-temperature limit, the system features a single unstable small black hole state. In the high-temperature limit, the system supports two unstable small black hole states along with one stable large black hole state. As the inverse temperature increases, the system shifts to a state with only one unstable small black hole. For temperatures within the range $\beta_{c2}=23.2166<\beta<\beta_{c3}=23.2196$, the system contains two unstable small black hole branches and one stable large black hole branch, yielding a total topological number of W=1W=-1. Notably, throughout all values $(\beta)$, the topological number remains consistently 1-1, and the system’s topological classification matches that of the $W^{1-}$ classes. As a result, we suggest that this system belongs to a new subclass, $\ddot{W}^{1-}$, within the topological classification framework. Furthermore, Table 3 provides a summary of the asymptotic behavior of the vector field near the boundaries $I_{i}$.
@@
In this analysis, we explore the asymptotic behavior of the newly discovered topological subclass. Initially, the system consists of at least three black hole states: with an increasing event horizon radius, the system transitions through three phases: two unstable states $(\omega=-1)$ and one stable state $(\omega=1)$. In these phases, the heat capacities are negative for the first two states and positive for the final, stable state. Subsequent states appear in pairs, where each pair includes one state with a winding number of 1-1 and one with +1+1. The arrangement of winding numbers does not follow a simple alternating pattern but instead begins with [,+][-,+]. This ensures that the innermost black hole is associated with $\omega=-1$ (an unstable state), and the outermost one with $\omega=+1$ (a stable state), forming a topological classification represented by [,+][-,+].
@@ IV Conclusions
Table 3: New topological subclasses: The orientation of the $\phi^{r_{h}}$ arrows and the associated topological numbers across the four segments.
Topological (sub)classes $I_{1}$ $I_{2}$ $I_{3}$ $I_{4}$ $W$
$\ddot{W}^{1-}$ $\leftarrow$ $\uparrow$ $\rightarrow$ $\downarrow$ -1
@@ Acknowledgments
We are grateful to Dr. Di Wu for useful discussions on the topological subclass. We also acknowledge the anonymous referees for their valuable comments on improving our paper. This work is supported by the National Natural Science Foundation of China (NSFC) under Grants No. 12265007, 12405081, 11865018. by the Doctoral Foundation of Zunyi Normal University of China (BS [2022] 07). Also, the research of H.H. was supported by the Q-CAYLE project, funded by the European Union-Next Generation UE/MICIU/Plan de Recuperacion, Transformacion y Resiliencia/Junta de Castilla y Leon (PRTRC17.11), and also by project PID2023-148409NB-I00, funded by MICIU/AEI/10.13039/501100011033. Financial support of the Department of Education of the Junta de Castilla y Leon and FEDER Funds is also gratefully acknowledged (Reference: CLU-2023-1-05). Additionally, H. H. is grateful to Excellence project FoS UHK 2203/2025-2026 for the financial support.